Fermi National Laboratory
 

Accelerator Update for January 5, 2007 - January 8, 2007

 

Friday January 5

The day shift began with Operations monitoring store 5162 (60.62E30) and a Recycler stash (216.9E10), with the Antiproton Source (Pbar) stacking (7.4E10), with Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE off due to absorber work, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

At 8:34 AM, the NuMI beam permit tripped off due to a beam position problem for the first of four times in two hours. It reset each time.

At 9:04 AM, a Pbar Accumulator PLC (A:PLC11) came into alarm. Operations contacted an expert who investigated and reported that the problem was in the tunnel. This will slow stacking until it is fixed.

A NuMI expert began a horn study at 9:06 AM.

At 10:10 AM, Operations stashed the Pbar antiprotons into the Recycler and then prepared Pbar for access. Pbar experts will repair the PLC. MI will take advantage of the access and repair a LCW leak on the MI-60 upper power supply.

At 11:38 AM, MI experts reported that the LCW leak was from the power supply’s choke and would take more than a day to replace. This replacement was put off and experts bypassed this power supply.

Pbar came out of access at 11:42 AM. Operations resumed sending beam to NuMI. Stacking resumed at 12:05 PM.

At 12:23 PM, Operations reported problems with the Pbar Debuncher cooling bands. Operations contacted an expert who said it was probably due to the new PLC and that he’d check it out.

At 1:04 PM, NuMI tripped off due to beam position problems for the first of three times in about three hours.

A Pbar expert completed his PLC work. Normal stacking (19.68mA/hr) resumed at 1:28 PM.

A MI expert began a 1.5-hour multi batch injection study at 3:14 PM.

Pbar experts began a stacktail heating study around 3:20 PM.

At 7:32 PM, a Booster vacuum pump tripped off and caused two RF stations (BRF13 & 14) to trip off. Everything reset okay.

Operations terminated store 5162 at 8:01 PM.

Operations prepared D0 for access at 8:20 PM. They were out of access by 10 PM.

The Recycler completed its last of three stashes for the evening shift at 10:16 PM.

Operations put the Tevatron (TeV) through a dry squeeze at 10:44 PM.

Operations put the TeV into shot setup at 11:12 PM.

Saturday January 6

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (352.0E10), with Pbar stacking on hold for shot setup, and with NuMI taking beam.

Operations established store 5164 at 1:35 AM with an initial luminosity of 240.28E30.

At 1:56 AM, a MI-60 water pump for a RF cavity, tripped off. The pump would not reset. The Duty Mechanic turned on the backup pump. This trip took down several Recycler RF stations. Operators eventually got these stations back online.

At 3:35 AM, the MI power supply interlock loop tripped off for the first of two times in an hour on a MI-50 upper power supply volts-to-ground fault. It reset both times.

The Recycler completed a stash at 7:53 AM.

At 8:18 AM, Operations reported that a Pbar Debuncher RF station (DRF1-2) had tripped off after the stash. It reset.

The Recycler completed a stash at 10:25 AM.

At 10:39 AM, DRF1-2 again tripped off after the stash. It reset.

The Recycler completed a stash at 1:01 PM. DRF1-2 again tripped off after the stash.

Operations reported at 8:24 PM that they had completed three more Recycler stashes without a DRF1-2 trip.

Sunday January 7

The midnight shift began with Operations monitoring store 5164 and the stash, with Pbar stacking, and with NuMI taking beam.

Operations terminated store 5164 at 2 AM and then put the TeV into shot setup.

Operations established store 5165 at 4:33 AM with an initial luminosity of 241.02E30.

The Recycler completed its last of two stashes for the midnight shift at 5:52 AM.

Operations reported at 6:21 PM that the Recycler had completed its fourth stash for the day.

At 10:36 PM, MI-10 suffered a vacuum burst. Everything reset okay.

Operations reported at 11:49 PM that the Recycler had completed two more stashes during the evening shift.

Monday January 8

The midnight shift began with Operations monitoring store 5165 and the stash, with Pbar stacking, and with NuMI taking beam.

The Recycler completed its last of two stashes for the midnight shift at 5:26 AM.

At 5:34 AM, Operations reported that a TeV RF station (TRF8) tripped off without an indication of a problem. Operators left the station off.

Operations terminated store 5165 at 8 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow MI and Booster experts to conduct studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/09/2007    email Fermilab